# Normalizing Flows for Gaussian Process Regression under Hierarchical Shrinkage Priors

### v1

@@ Abstract
Gaussian Process Regression (GPR) is a powerful tool for nonparametric regression, but its fully Bayesian application in high-dimensional settings is hindered by two primary challenges: the computational burden (exacerbated by fully Bayesian inference) and the difficulty of variable selection. This paper introduces a novel methodology that combines hierarchical global-local shrinkage priors with normalizing flows to address these challenges. The hierarchical triple gamma prior offers a principled framework for inducing sparsity in high-dimensional GPR, effectively excluding irrelevant covariates while preserving interpretability and flexibility in model size. Normalizing flows are employed within a variational inference framework to approximate the posterior distribution of hyperparameters, capturing complex dependencies while ensuring computational scalability. Simulation studies demonstrate the efficacy of the proposed approach, outperforming traditional maximum likelihood estimation and mean-field variational methods, particularly in high-sparsity and high-dimensional settings. The results highlight the robustness and flexibility of hierarchical shrinkage priors and the computational efficiency of normalizing flows for Bayesian GPR. This work provides a scalable and interpretable solution for high-dimensional regression, with implications for sparse modeling and posterior approximation in broader Bayesian contexts.
62G08,
62F15,
62H12,
60G15,
Variational inference,
Variable selection,
Covariance shrinkage,
Triple gamma shrinkage prior,
keywords:
[class=MSC]
keywords:
volume: TBAissue: TBA
[…]
@@ 1 Introduction
Gaussian process regression (GPR) is a widely used tool in nonparametric regression, valued for its theoretical rigor—achieving near-minimax rates of convergence to the true underlying function (van der Vaart and van Zanten, 2009)—all while retaining interpretability. However, two major challenges arise in practice. First, selecting the hyperparameters that govern the covariance structure becomes increasingly complex as the number of predictors grows, requiring variable selection to distinguish influential covariates from noise. Second, estimating GPR models, particularly within a fully Bayesian framework, is computationally demanding. The cost of evaluating the likelihood scales cubically with the number of observations, making standard approaches like Markov chain Monte Carlo (MCMC) prohibitively expensive for large datasets.
@@
Several methods have been proposed for variable selection in GPR. The most well-known is automatic relevance determination (ARD) by Neal (1996), which identifies covariate importance through maximum-likelihood hyperparameter estimates but lacks explicit sparsity. Penalized approaches, such as those by Yan and Qi (2010) and Yi et al. (2011), address this by enforcing sparsity directly, for example, through the $\ell_{1}$ penalty. Alternatively, sensitivity-based methods (e.g., Piironen and Vehtari, 2016; Paananen et al., 2019) evaluate the predictive changes when covariates are removed, while dimension-reduction techniques (e.g., Park et al., 2022; Tripathy et al., 2016; Liu and Guillas, 2017) focus on finding lower-dimensional representations of the covariates, often at the expense of interpretability.
@@
Despite being couched in Bayesian terminology, all previously mentionend methods for variable selection in GPR are not fully Bayesian, as they yield only point estimates for hyperparameters and lack the full uncertainty quantification inherent to Bayesian analysis. This limitation can result in overconfidence in predictions and reduced out-of-sample performance. Fully Bayesian approaches, while less common, aim to address this. For example, Chen and Wang (2010) and Savitsky et al. (2011) use spike-and-slab priors (Mitchell and Beauchamp, 1988; George and McCulloch, 1993, 1997) to induce sparsity, providing posterior inclusion probabilities for each covariate. However, these methods become computationally expensive in higher dimensions. Others, such as Vo and Pati (2017), adopt computationally tractable alternatives like the horseshoe prior (Carvalho et al., 2010) to induce sparsity in additive GPR frameworks. More recently, Tang et al. (2024) applied cumulative shrinkage priors (Legramanti et al., 2020) using the Karhunen-Loève decomposition (Alexanderian, 2015), which embeds a hierarchy of effects but sacrifices interpretability.
@@
While theoretically more attractive, fully Bayesian approaches face a significant computational hurdle, particularly in high-dimensional GPR settings, due to the cubic scaling of the likelihood evaluation with the number of observations. Sampling-based methods such as MCMC exacerbate this issue, as they require repeated likelihood evaluations. For example, single-move Metropolis-Hastings updates each parameter sequentially, making high-dimensional applications infeasible. Even more efficient samplers, like Hamiltonian Monte Carlo (HMC), can remain prohibitive due to the computational cost of evaluating gradients at each leapfrog step. Variational inference (VI) offers a potential solution by reducing computational demands, but many widely used VI implementations rely on mean-field approximations. This reliance limits expressivity, making it challenging to accurately capture complex posterior distributions and often leading to suboptimal results. These computational challenges constrain the applicability of fully Bayesian GPR to small datasets.
@@
In this paper, I tackle the two aforementioned problems simultaneously. Recent insights from the hierarchical global-local shrinkage literature are introduced into the GPR framework, providing an interpretable and flexible approach to variable selection. Additionally, normalizing flows (Rezende and Mohamed, 2015) are employed within a variational inference framework to flexibly approximate the posterior distribution of the hyperparameters, capturing dependencies in the posterior while ensuring scalability and maintaining proper uncertainty quantification. Together, these advances enable a principled Bayesian approach to variable selection, yielding interpretable models that remain computationally tractable in higher dimensions.
@@
The remainder of this paper is structured as follows. Section 2 introduces hierarchical global-local shrinkage priors for GPR. Section 3 provides an overview of normalizing flows and how they can be used to approximate the posterior distribution of the hyperparameters. Section 4 presents a simulation study to assess the performance of the proposed method. Section 5 applies the method to a real data set. Finally, Section 6 concludes.
@@ 2 Hierarchical Shrinkage in Gaussian Process Regression
$y_{i}=f(\mathbf{x}_{i})+\epsilon_{i},$ (1)
where $f$ is an unknown mean regression function to be estimated from the data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, which in turn consists of the covariates $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the responses $y_{i}\in\mathbb{R}$, and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the noise term. The regression function $f$ is placed under a Gaussian Process (GP) prior with a zero mean function (a choice that can be relaxed if needed). The GP prior is defined as
$f\mid\boldsymbol{\zeta},\mathbf{x}\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime};\boldsymbol{\zeta})),$ (2)
where $k(\mathbf{x},\mathbf{x}^{\prime};\boldsymbol{\zeta})$ is a positive definite kernel function that depends on the hyperparameters $\boldsymbol{\zeta}$. Under this specification, marginalizing over the latent function $f$ yields the prior joint distribution of the response vector $\mathbf{y}$, incorporating both the covariance structure and noise term:
@@
where the covariance matrix $\mathbf{K}(\mathbf{x};\boldsymbol{\zeta})$ is defined elementwise as $\mathbf{K}(\mathbf{x};\boldsymbol{\zeta})_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j};\boldsymbol{\zeta})$ (Rasmussen and Williams, 2005).
Many choices are feasible for the covariance function $k(\mathbf{x},\mathbf{x}^{\prime};\boldsymbol{\zeta})$. Examples include the squared exponential kernel, the Matérn kernel, the rational quadratic kernel, or the periodic kernel. While these differ in characteristics like implied smoothness, correlation between ”distant” observations, or periodicity, many of them — when used in the context of regression — have at their core an anisotropic distance function of the form
$\delta(\mathbf{x},\mathbf{x}^{\prime};\boldsymbol{\theta})=\sqrt{\sum_{j=1}^{d}(x_{j}-x_{j}^{\prime})^{2}\theta_{j}},$ (4)
@@
$k_{SE}(\mathbf{x},\mathbf{x}^{\prime};\tau,\boldsymbol{\theta})=\frac{1}{\tau}\exp\left(-\frac{1}{2}\delta(\mathbf{x},\mathbf{x}^{\prime};\boldsymbol{\theta})^{2}\right).$ (5)
where $\tau$ determines the overall variance of the Gaussian process. The hyperparameters $\boldsymbol{\theta}$ dictate how each covariate contributes to the distance function, directly influencing the covariance structure of the Gaussian process. As $\theta_{j}$ becomes small, the $j$-th covariate contributes less to the distance function, resulting in reduced influence on the covariance structure. In the limit, when $\theta_{j}=0$, the $j$-th covariate no longer affects the covariance structure and can be considered irrelevant. This effect is clearly illustrated in Figure 1, which shows realizations from a Gaussian process with varying $\theta_{j}$ values. Smaller values of $\theta_{j}$ result in smoother functions, while $\theta_{j}=0$ leads to a constant function, where the covariate has no more influence on the outcome. Notably, the covariance structure is governed by only a handful of hyperparameters, which simplifies the estimation process discussed in Section 3. Despite this simplicity, the Gaussian process retains high flexibility, making it suitable for modeling complex relationships.
Figure 1: Samples from a Gaussian process with a squared exponential kernel, with varying values of $\theta_{j}$. As $\theta_{j}$ becomes smaller, the draws become smooother, with the extreme case of $\theta_{j}=0$ leading to a constant function.
When $d$ is large, i.e., when one is faced with many potential predictors, a mechanism is required to remove irrelevant covariates from the model. In high-dimensional settings, it is often the case that many covariates are irrelevant, contributing noise rather than predictive value. Removing these covariates improves both interpretability and statistical efficiency while safeguarding predictive performance. In the more traditional linear regression setting, this is a well-studied problem, with a popular Bayesian approach being global-local shrinkage priors (Polson and Scott, 2011):
@@
where $\beta_{j}$ is the regression coefficient for the $j$-th covariate, $\lambda_{j}$ is the local shrinkage parameter, and $\tau$ is the global shrinkage parameter. The idea behind these priors is that $\tau$ pulls all $\beta_{j}$ towards zero, while local parameters $\lambda_{j}$ allow select $\beta_{j}$ to remain non-zero if supported by the data. Such priors are hierarchical in nature, with the global parameter $\tau$ inducing a joint distribution that shares information across coefficients and enables adaptation to varying sparsity levels. Various choices of $\pi_{\lambda}(\lambda_{j})$ and $\pi_{\tau}(\tau)$ have been studied in the literature, such as the normal-gamma prior (Griffin and Brown, 2010), the Bayesian Lasso (Park and Casella, 2008), the horseshoe prior (Carvalho et al., 2009) or the triple gamma prior (Cadonna et al., 2020), to name a few.
In their default specification, global-local shrinkage priors are not directly applicable to the GPR framework, as they are defined for regression coefficients with support on $\mathbb{R}$. In contrast, the parameters that govern the covariance structure in GPR, $\theta_{1},\dots,\theta_{d}$, must remain positive to ensure a valid distance function and meaningful covariance structure. To address this, this work adopts the approach of Frühwirth-Schnatter and Wagner (2010), which is well-established in the time-varying parameter literature, and sets $\theta_{j}=\beta_{j}^{2}$. This transformation replaces the normal distribution at the lowest level of the hierarchy with a gamma distribution:
@@
which defines a class of positive shrinkage priors, while retaining the global-local structure, enabling shared shrinkage across parameters via $\tau$ and flexibility for individual $\theta_{j}$ through $\lambda_{j}$.
This work specifically considers the hierarchical triple gamma prior of Cadonna et al. (2020) for $\theta_{j}$, which is well-suited for high-dimensional GPR due to its flexibility, generality, and computational tractability. While the triple gamma prior is traditionally represented as a compound distribution of three gamma random variables, its global-local form closest to (7) is given by
$\theta_{j}\mid\tau,\lambda_{j}\sim\mathcal{G}\left(\frac{1}{2},\frac{1}{2\tau\lambda_{j}}\right),\quad\lambda_{j}\mid a,c\sim F\left(2a,2c\right),\quad\tau\mid a,c\sim F\left(2c,2a\right).$ (8)
The triple gamma prior offers three key advantages. First, it is a generic choice of shrinkage prior, encompassing many well-known priors as special cases (e.g., the other global-local shrinkage priors discussed previously). Second, it is highly flexible, with the parameters $a$ and $c$ modifying the pole and tail behavior, respectively. Specifically, the pole around the origin becomes more pronounced as $a$ decreases, while the tails of the distribution become heavier as $c$ decreases. The pronounced pole effectively squelches noise, while the heavy tails allow signals to filter through. Finally, and importantly for the estimation procedure presented in Section 3, the marginal density is available in closed form:
$f(x;a,c,\tau)=\frac{\Gamma(c+{\frac{1}{2}})}{\sqrt{2\pi\kappa x}B(a,c)}U\left(c+{\frac{1}{2}},\frac{3}{2}-a,\frac{x}{2\kappa}\right),$ (9)
where $\kappa=\frac{2c}{\tau a}$ and $U\left(a,b,z\right)$ is the confluent hypergeometric function of the second kind:
[…]
The closed-form marginal density simplifies the estimation procedure by reducing reliance on the full hierarchical structure of the prior, keeping the number of hyperparameters to estimate low. This computational efficiency, combined with the prior’s flexibility and generality, makes it a strong choice for high-dimensional regression tasks.

### v2

# Going With the Flow: Normalizing Flows for Gaussian Process Regression under Hierarchical Shrinkage Priors

@@ Abstract
Gaussian Process Regression (GPR) is a powerful tool for nonparametric regression, but its application in a fully Bayesian fashion in high-dimensional settings is hindered by two primary challenges: the difficulty of variable selection and the computational burden, which is particularly acute in fully Bayesian inference. This paper introduces a novel methodology that combines hierarchical global-local shrinkage priors with normalizing flows to address these challenges. The hierarchical triple gamma prior offers a principled framework for inducing sparsity in high-dimensional GPR, effectively excluding irrelevant covariates while preserving interpretability and flexibility. Normalizing flows are employed within a variational inference framework to approximate the posterior distribution of parameters, capturing complex dependencies while ensuring computational scalability. Simulation studies demonstrate the efficacy of the proposed approach, outperforming traditional maximum likelihood estimation and mean-field variational methods, particularly in high-sparsity and high-dimensional settings. This is also borne out in an application to binding affinity ($\text{pIC}_{50}$) measurements for small molecules targeting $\beta$-secretase-1 (BACE-1). The results highlight the robustness and flexibility of hierarchical shrinkage priors and the computational efficiency of normalizing flows for Bayesian GPR. This work provides a scalable and interpretable solution for high-dimensional nonparametric regression, with implications for sparse modeling and posterior approximation in broader Bayesian contexts.
Keywords: Variational inference, Variable selection, Covariance shrinkage, Triple gamma shrinkage prior
@@ 1 Introduction
Gaussian process regression (GPR) is a widely used tool in nonparametric regression, valued for its theoretical rigor. It achieves near-minimax rates of convergence to a true underlying function (vanderVaart2009), all while retaining a higher degree of interpretability than fully black box methods. However, two major challenges arise in practice. First, selecting the hyperparameters that govern the covariance structure becomes increasingly complex as the number of predictors grows, requiring variable selection to distinguish influential covariates from noise. Second, estimating GPR models, particularly within a fully Bayesian framework, is computationally demanding. The cost of evaluating the marginal likelihood scales cubically with the number of observations, making standard approaches like Markov chain Monte Carlo (MCMC) prohibitively expensive for even moderately sized datasets.
@@
Several methods have been proposed for variable selection in GPR. The most well-known is automatic relevance determination (ARD) by neal1996bayesian, which identifies covariate importance through maximum-likelihood estimates of the parameters but does not impose sparsity explicitly. Penalized approaches, such as those by yan2010sparse and yi2011penalized, address this by enforcing sparsity directly through, for example, the $\ell_{1}$ penalty. Alternatively, sensitivity-based methods (e.g., piironen2016projection, paananen2019variable) evaluate the predictive changes when covariates are removed, while dimension-reduction techniques (e.g., park2022variable, tripathy2016gaussian, liu2017dimension) focus on finding lower-dimensional representations of the covariates, often at the expense of interpretability.
@@
Despite being couched in Bayesian terminology, all previously mentioned methods for variable selection in GPR are not fully Bayesian, as they yield only point estimates for hyperparameters and lack the full uncertainty quantification inherent to Bayesian analysis. This limitation can result in overconfidence in predictions and reduced out-of-sample performance. Fully Bayesian approaches, while less common, aim to address this. For example, chen2010bayesian and savitsky2011variable use spike-and-slab priors (mitchell1988bayesian, george1993variable, george1997approaches) to induce sparsity, providing posterior inclusion probabilities for each covariate. However, these methods become computationally expensive in higher dimensions. Others, such as vo2017sparse, adopt computationally tractable alternatives like the horseshoe prior (carvalho2010horseshoe) to induce sparsity in additive GPR frameworks. More recently, tang2023hierarchical applied cumulative shrinkage priors (legramanti2020bayesian) using the Karhunen-Loève decomposition (alexanderian2015brief), which embeds a hierarchy of effects but sacrifices interpretability.
@@
While theoretically more attractive, fully Bayesian approaches such as MCMC often particularly suffer from the high computational cost of evaluating the marginal likelihood. Consider single-move Metropolis-Hastings updates. Each parameter is updated sequentially, with each update requiring an evaluation of the marginal likelihood, making high-dimensional applications infeasible. Even more efficient samplers, like Hamiltonian Monte Carlo (HMC), can remain prohibitive due to the computational cost of evaluating gradients at each leapfrog step. Variational inference offers a potential solution by reducing computational demands, but many widely used implementations rely on mean-field approximations. This choice of variational family limits expressivity, making it virtually impossible to accurately capture complex posterior distributions. Taken together, these computational challenges constrain the applicability of fully Bayesian GPR to small datasets.
@@
This paper tackles the two aforementioned problems simultaneously. Recent insights from the hierarchical global-local shrinkage literature are introduced into the GPR framework, providing an interpretable and flexible approach to variable selection. Additionally, normalizing flows (rezende2015variational) are employed within a variational inference framework to flexibly approximate the posterior distribution of the hyperparameters, capturing dependencies in the posterior while ensuring scalability and maintaining proper uncertainty quantification. Together, these advances enable a principled Bayesian approach to variable selection in GPR, yielding flexible, interpretable models that remain computationally tractable in higher dimensions. To facilitate practical application, all methods introduced in this paper are implemented in the R package shrinkGPR (shrinkGPR), which is available on CRAN.
@@
The remainder of this paper is structured as follows. Section 2 introduces hierarchical global-local shrinkage priors for GPR. Section 3 provides an overview of normalizing flows and how they can be used to approximate the posterior distribution of the parameters. Section 4 presents a simulation study to assess the performance of the proposed method. Section 5 applies the method to a dataset containing binding affinity ($\text{pIC}_{50}$) measurements for molecules targeting $\beta$-secretase-1 (BACE-1), a key enzyme implicated in Alzheimer’s disease. Finally, Section 6 concludes.
@@ 2 Hierarchical Shrinkage in Gaussian Process Regression
[…]
where $f$ is an unknown mean regression function to be estimated from the data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, which in turn consists of the covariates $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the responses $y_{i}\in\mathbb{R}$. $\epsilon_{i}$ is the noise term, which is independent of the covariates. The regression function $f$ is modeled as a random function drawn from a Gaussian process (GP) prior with a zero mean function (a choice that can be relaxed if needed). A GP is a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$:
$f\mid\boldsymbol{\zeta}\sim\mathcal{GP}(0,k(\cdot,\cdot;\boldsymbol{\zeta})),$
where $k$ is a positive definite kernel function depending on hyperparameters $\boldsymbol{\zeta}$. This prior implies that, for any finite set of covariates $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, the vector of function evaluations $\mathbf{f}=(f(\mathbf{x}_{1}),\dots,f(\mathbf{x}_{N}))^{\top}$ follows a multivariate normal distribution. Integrating out the latent function then yields a marginal distribution for the observed response vector:
@@
where the covariance matrix $\mathbf{K}(\mathbf{x};\boldsymbol{\zeta})$ is defined elementwise as $\mathbf{K}(\mathbf{x};\boldsymbol{\zeta})_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j};\boldsymbol{\zeta})$ (rasmussen2005gaussian), and $\mathbf{I}$ is the $N\times N$ identity matrix.
Many choices are feasible for the covariance function $k(\cdot,\cdot;\boldsymbol{\zeta})$. Examples include the squared exponential kernel, the Matérn kernel, the rational quadratic kernel, or the periodic kernel. While these differ in characteristics like implied smoothness, correlation between "distant" observations, or periodicity, many of them (when used in the context of regression) have at their core an anisotropic distance function of the form
$\delta(\mathbf{z},\mathbf{z}^{\prime};\boldsymbol{\theta})=\sqrt{\sum_{j=1}^{d}(z_{j}-z_{j}^{\prime})^{2}\theta_{j}},$ (3)
@@
$k_{\text{SE}}(\mathbf{z},\mathbf{z}^{\prime};\tau,\boldsymbol{\theta})=\frac{1}{\tau}\exp\left(-\frac{1}{2}\delta(\mathbf{z},\mathbf{z}^{\prime};\boldsymbol{\theta})^{2}\right),$ (4)
where $\tau$ determines the overall variance of the GP. In this case, the hyperparameters that control the covariance structure are $\boldsymbol{\zeta}=\{\boldsymbol{\theta},\tau\}$. The parameters contained in $\boldsymbol{\theta}$ dictate how each covariate contributes to the distance function, in turn directly influencing the covariance structure of the GP. As $\theta_{j}$ becomes small, the $j$-th covariate contributes less to the distance function, resulting in reduced influence on the covariance structure. In the limit, when $\theta_{j}=0$, the $j$-th covariate no longer affects the covariance structure and can be considered irrelevant. This effect is clearly illustrated in Figure 1, which shows realizations from a GP with varying $\theta_{j}$ values. Smaller values of $\theta_{j}$ result in smoother functions, while $\theta_{j}=0$ leads to a constant function, where the covariate has no more influence on the outcome. Notably, the covariance structure is governed by only a handful of hyperparameters, which simplifies the estimation process discussed in Section 3. Despite this simplicity, the GP retains high flexibility, making it suitable for modeling complex relationships.
Figure 1: Samples from a GP with a squared exponential kernel, with varying values of $\theta_{j}$. As $\theta_{j}$ becomes smaller, the draws become smooother, with the extreme case of $\theta_{j}=0$ leading to a constant function.
When $d$ is large, i.e., when one is faced with many potential predictors, a mechanism is required to remove irrelevant covariates from the model. In such high-dimensional settings, a reasonable assumption is that many covariates are irrelevant, contributing noise rather than predictive value. In the more traditional linear regression setting, this is a well-studied problem, with a popular Bayesian approach being global-local shrinkage priors (polson2011shrink):
@@
where $\beta_{j}$ is the regression coefficient for the $j$-th covariate, $\lambda_{j}$ is the local shrinkage parameter, and $\tau$ is the global shrinkage parameter. The idea behind these priors is that $\tau$ pulls all $\beta_{j}$ towards zero, while local parameters $\lambda_{j}$ allow select $\beta_{j}$ to remain non-zero if supported by the data. Such priors are hierarchical in nature, with the global parameter $\tau$ inducing a joint distribution that shares information across coefficients and enables adaptation to varying sparsity levels. Various choices of $\pi_{\lambda}(\lambda_{j})$ and $\pi_{\tau}(\tau)$ have been studied in the literature, such as the normal-gamma prior (gri-bro:inf), the Bayesian Lasso (par-cas:bay), the horseshoe prior (carvalho2010horseshoe) or the triple gamma prior (cad-etal:tri), to name a few.
In their default specification, global-local shrinkage priors are not directly applicable to the GPR framework, as they are defined for regression coefficients with support on $\mathbb{R}$. In contrast, the parameters that govern the covariance structure in GPR, $\theta_{1},\dots,\theta_{d}$, must remain positive to ensure a valid distance function and meaningful covariance structure. To address this, this work adopts the approach of fruhwirth2010stochastic, which is well-established in the time-varying parameter literature, and sets $\theta_{j}=\beta_{j}^{2}$. This transformation replaces the normal distribution at the lowest level of the hierarchy with a gamma distribution:
@@
which defines a class of shrinkage priors with support on $\mathbb{R}^{+}$, while retaining the global-local structure, enabling shared shrinkage across parameters via $\tau$ and flexibility for individual $\theta_{j}$ through $\lambda_{j}$.
This work specifically considers the hierarchical triple gamma prior of cad-etal:tri for $\theta_{j}$, which is well-suited for high-dimensional GPR due to its flexibility, generality, and computational tractability. While the triple gamma prior is traditionally represented as a compound distribution of three gamma random variables (hence the name), its global-local form closest to (6) is given by
$\theta_{j}\mid\tau,\lambda_{j}\sim\mathcal{G}\left(\frac{1}{2},\frac{1}{2\tau\lambda_{j}}\right),\quad\lambda_{j}\mid a,c\sim F\left(2a,2c\right),\quad\tau\mid a,c\sim F\left(2c,2a\right),$ (7)
where $F(d_{1},d_{2})$ is the F distribution with degrees of freedom $d_{1}$ and $d_{2}$. The triple gamma prior offers three key advantages. First, it is a generic choice of shrinkage prior, encompassing many well-known priors as special cases (e.g., the other global-local shrinkage priors discussed previously). Second, it is highly flexible, with the parameters $a$ and $c$ modifying the pole and tail behavior, respectively. Specifically, the pole around the origin becomes more pronounced as $a$ decreases, while the tails of the distribution become heavier as $c$ decreases. The pronounced pole effectively squelches noise, while the heavy tails allow signals to filter through. Finally, and importantly for the estimation procedure presented in Section 3, the prior on $\theta_{j}$ under the triple gamma marginalised with respect to $\lambda_{j}$ is available in closed form:
$f(\theta_{j};a,c,\tau)=\frac{\Gamma(c+{\frac{1}{2}})}{\sqrt{2\pi\kappa\theta_{j}}B(a,c)}U\left(c+{\frac{1}{2}},\frac{3}{2}-a,\frac{\theta_{j}}{2\kappa}\right),$ (8)
where $\kappa=\tau\frac{c}{a}$ and $U\left(a,b,z\right)$ is the confluent hypergeometric function of the second kind:
$U\left(a,b,z\right)=\frac{1}{\Gamma(a)}\int_{0}^{\infty}e^{-zt}t^{a-1}(1+t)^{b-a-1}dt.$
The closed-form marginal density simplifies the estimation procedure by collapsing the full hierarchical structure of the prior, keeping the number of hyperparameters to estimate per individual $\theta_{j}$ low. This computational efficiency, combined with the prior’s flexibility and generality, makes it a strong choice for the GPR framework presented here.